How many molecules of oxygen are in one breath?

an animal takes a breath with a volume of 5.80 L at standard temperature and pressure. If the air contains 20.9% oxygen.

................O2 molecules


1
Expert's answer
2021-03-08T06:12:54-0500

V(O2)=5.80*0.209=1.212(L)

PV=nRT

n=PV/RT

n=1*1.212/0.082*273.15 =0.054(mol)

N=n*Na=0.054*6.02•1023=3.25•1022

Answer: 3.25•1022 O2 molecules.